OpenStudy (geekfromthefutur):

A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (geekfromthefutur):

.@terenzreignz

OpenStudy (unklerhaukus):

\[[\text{kB}]=2^{10}[\text{B}]\qquad[\text{MB}]=2^{20}[\text{B}]\] \[\frac{[\text{MB}]}{[\text{kB}]}=\frac{2^{20}[\text{B}]}{2^{10}[\text{B}]}=\]

OpenStudy (unklerhaukus):

Use \[\frac{x^n}{x^m}=x^{n-m}\]

OpenStudy (unklerhaukus):

what did you get?